OpenStudy (anonymous):

Let h and g be two functions defined by H(x)=2x^2-5 G(x)=3x-8 For any real number x Find h/g(-3) and then find all values that are not in the domain of h/g. If more than one value separate with commas.

OpenStudy (anonymous):

h/g = h(x)/g(x) = (2x^2-5)/(3x-8) h/g is undefined when the denominator 3x-8 is equal to zero. this value of x will not be in the domain of the function

OpenStudy (anonymous):

Thank you